1)Suppose that the population distribution of the gripping strengths of industrial workers is known to have a mean of 110 and standard deviation of 10. For a random sample of 75 workers, what is the probability that the sample mean gripping strength will be

a)Between 109 and 112? 

b)Greater than112?

2)On the average, five smokers pass a certain street corners every ten minutes, what is the probability that during a given 10minutes the number of smokers passing will be 

A)6 or fewer 

B)7 or more 

C)Exactly 8


1
Expert's answer
2021-05-13T01:20:11-0400

1)

"\mu = 110 \\

\sigma = 10 \\

n = 75 \\"

a)

"P(109<X<112) = P(X<112)-P(X<109) \\

= P(Z< \frac{112-110}{10/ \sqrt{75}}) -P(Z< \frac{109-110}{10/ \sqrt{75}}) \\

= P(Z<1.733)-P(Z< -0.866) \\

= 0.9584 -0.1932 \\

= 0.7652"

b)

"P(X>112) = 1-P(X<112) \\

= 1 -P(Z< \frac{112-110}{10/ \sqrt{75}}) \\

= 1 -P(Z<1.733) \\

= 1 -0.9584 \\

= 0.0416"

2) Assuming that the number of smokers follows Poisson distribution, if X is a Poisson random variable with parameter λ then

"P(X=x) = \frac{λ^x \times e^{-λ}}{x!} \\

λ = 5"

A) P(X≤6) = P(X=0) + P(X=1) + P(X=2) + P(X=3) + P(X=4) + P(X=5) + P(X=6)

"P(X=0) = \frac{5^0 \times e^{-5}}{0!} \\

= \frac{0.006737}{1} \\

= 0.006737 \\

P(X=1) = \frac{5^1 \times e^{-5}}{1!} \\

= \frac{0.03368}{1} \\

= 0.03368 \\

P(X=2) = \frac{5^2 \times e^{-5}}{2!} \\

= \frac{0.16844}{2} \\

= 0.08422 \\

P(X=3) = \frac{5^3 \times e^{-5}}{3!} \\

= \frac{0.84224}{6} \\

= 0.14037 \\

P(X=4) = \frac{5^4 \times e^{-5}}{4!} \\

= \frac{4.21121}{24} \\

= 0.17546 \\

P(X=5) = \frac{5^5 \times e^{-5}}{5!} \\

= \frac{21.056}{120} \\

= 0.17546 \\

P(X=6) = \frac{5^6 \times e^{-5}}{6!} \\

= \frac{105.28}{720} \\

= 0.14622 \\

P(X≤6) = P(X=0) + P(X=1) + P(X=2) + P(X=3) + P(X=4) + P(X=5) + P(X=6) \\

P(X≤6) = 0.006737 + 0.03368 + 0.08422 + 0.14037 + 0.17546 + 0.17546 + 0.14622 \\

= 0.762447"

B) P(X≥7) = 1 -P(X<7)

P(X<7) = P(≤6)

"P(X≤6) = 0.762447 \\

P(X≥7) = 1 -0.762447 \\

= 0.237553"

C) the probability that, during a given 10-minute period, the number of smokers passing the street corner will be eight is:

"P(X=8) = \frac{5^8 \times e^{-5}}{8!} \\

= \frac{2632.01}{40320} \\

= 0.065"